OpenStudy (anonymous):

Marla drives her car to school each day. If it takes her 12 minutes to drive one way, how many hours a week does Marla spend driving? (Express the answer in decimal form.)

OpenStudy (anonymous):

one day is \(2\times 12\) miles because she goes there and back if she does this 5 days a week it will be \(5\times 2\times 12\)

OpenStudy (anonymous):

@satellite73 i get 120

OpenStudy (anonymous):

me too then you have to convert 120 minutes in to hours, which should be easy enough, since there are 60 minutes in an hour